# jina-embeddings-v3: Multilingual Embeddings With Task LoRA

### v2

@@ Abstract
We introduce jina-embeddings-v3, a novel text embedding model with 570 million parameters, achieves state-of-the-art performance on multilingual data and long-context retrieval tasks, supporting context lengths of up to 8192 tokens. The model includes a set of task-specific Low-Rank Adaptation (LoRA) adapters to generate high-quality embeddings for query-document retrieval, clustering, classification, and text matching. Additionally, Matryoshka Representation Learning is integrated into the training process, allowing flexible truncation of embedding dimensions without compromising performance. Evaluation on the MTEB benchmark shows that jina-embeddings-v3 outperforms the latest proprietary embeddings from OpenAI and Cohere on English tasks, while achieving superior performance compared to multilingual-e5-large-instruct across all multilingual tasks.
@@ 4.1 Pre-Training
For long-context support, we first train for $100,000$ steps on text sequences of 512 tokens, followed by an additional $60,000$ steps using on text sequences of 512 tokens and a reduced batch size. The details are provided in Appendix A1.
@@ 4.2 Fine-Tuning for the Embedding Task
The training data consists of over one billion text pairs, drawn from more than 300 distinct sub-datasets, each representing specific domains in various languages. During training, the data loader constructs each batch by sampling a specific sub-dataset, ensuring that only text pairs from that dataset are included.
@@ 4.3 Training Task-Specific Adapters
Related work on embedding training Wang et al. (2022); Xiao et al. (2023) introduces an additional generic training phase following the pair-wise training phase. This phase incorporates high-quality data from various tasks to optimize model performance across a range of downstream use cases. In this stage, recent approaches use task-specific instructions to help the model distinguish between different tasks and domains, as discussed in Section 2.3.
However, this approach complicates model usage, as users must learn task-specific instructions (i.e., prompts) that align with the model’s behavior or "vibe." While this offers flexibility, it also makes the model’s behavior harder to predict. In contrast, we train five distinct LoRA adapters for four well-defined task types: classification, text matching, asymmetric retrieval (e.g. query-passage and query-document retrieval), and separation (e.g. clustering and reranking). These tasks are trained independently, with the base model’s weights kept frozen. For asymmetric retrieval, two adapters are trained jointly: one for queries and one for passages. During inference, users can select the appropriate adapter based on their downstream task and input role, ensuring optimal embeddings for their specific use case.
@@ 4.3.1 Classification Adapter
The classification adapter generates embeddings that are effective for training downstream classification models, such as logistic regression classifiers. To train the adapter, we employ the classification training method proposed for the Gecko embedding model Lee et al. (2024a). Specifically, we select datasets covering a range of common classification tasks, including sentiment analysis, intent classification, and article categorization.
@@
We employ an extended version of the InfoNCE loss $\mathcal{L}_{\text{triplet}}$ described in our previous work (CITATION HERE) to take into account these additional negative samples.
@@ 4.3.4 Failure Analysis for Asymmetric Retrieval
The effectiveness of this approach is evaluated in Section 5.4.
@@ 5.2 Performance on MTEB
Table 3 summarizes the performance of various multilingual text embedding models across different MTEB tasks, divided into English and multilingual sections. jina-embeddings-v3 performs competitively, particularly in monolingual English tasks, where it achieves the highest Classification Accuracy (CF) score of 82.58 and the top Sentence Similarity (STS) score of 85.8, demonstrating its robustness across both languages and tasks. Full evaluation results per task can be found in Appendix A3. When averaging across all tasks, jina-embeddings-v3 scores 65.60, outperforming models such as text-embedding-3-large, multilingual-e5-large-instruct, and Cohere-embed-multilingual-v3.0. This indicates that jina-embeddings-v3 maintains strong monolingual English performance while being trained on a wide variety of languages.
@@
For multilingual performance, Table 3 presents weighted averages across a wide selection of multilingual and cross-lingual MTEB tasks. The specific tasks, along with an explanation of which adapter was used for each task, are detailed in Appendix A7, A6, A8, A5, A9, and A4. Note that jina-embeddings-v2 refers to our bilingual model suite (jina-embeddings-v2-base-zh, jina-embeddings-v2-base-es, jina-embeddings-v2-base-de), which are evaluated only on Chinese, Spanish, and German monolingual and cross-lingual tasks, excluding all other tasks. We do not report scores for text-embedding-3-large and Cohere-embed-multilingual-v3.0 as these models were not evaluated on the full range of multilingual and cross-lingual MTEB tasks. However, our model outperforms multilingual-e5-large in all multilingual tasks except reranking and approaches the performance of multilingual-e5-large-instruct.
@@
Regarding the bilingual suite of models under jina-embeddings-v2, only tasks involving Chinese, Spanish, German, and English are reported in Table 3. Thus, the average score for jina-embeddings-v2* on the pair-classification tasks reflects performance on Chinese tasks alone. A complete report on pair-classification can be found in Appendix A6.

### v3

@@ Abstract
We introduce jina-embeddings-v3, a novel text embedding model with 570 million parameters, achieves state-of-the-art performance on multilingual data and long-context retrieval tasks, supporting context lengths of up to 8192 tokens. The model includes a set of task-specific Low-Rank Adaptation (LoRA) adapters to generate high-quality embeddings for query-document retrieval, clustering, classification, and text matching. Evaluation on the MTEB benchmark shows that jina-embeddings-v3 outperforms the latest proprietary embeddings from OpenAI and Cohere on English tasks, while achieving superior performance compared to multilingual-e5-large-instruct across all multilingual tasks. With a default output dimension of 1024, users can flexibly reduce the embedding dimensions to as low as 32 without compromising performance, enabled by Matryoshka Representation Learning.
@@ 4.1 Pre-Training
For long-context support, we first train for $100,000$ steps on text sequences that are truncated to 512 tokens, followed by an additional $60,000$ steps using on text sequences truncated to 8192 tokens and a reduced batch size. The details are provided in Appendix A1.
@@ 4.2 Fine-Tuning for the Embedding Task
The training data consists of over one billion text pairs, drawn from more than 300 distinct sub-datasets, each representing specific domains in various languages. During training, the data loader constructs each batch by sampling a specific sub-dataset, ensuring that only text pairs from that dataset and language are included.
@@ 4.3 Training Task-Specific Adapters
Related work on embedding training Wang et al. (2022); Xiao et al. (2023) introduces an additional generic training phase following the pair-wise training phase. This phase incorporates high-quality data from various task types to optimize model performance across a range of downstream use cases. In this stage, recent approaches use task-specific instructions to help the model distinguish between different tasks and domains, as discussed in Section 2.3.
However, this approach complicates the usage, as users must learn task-specific instructions (i.e., prompts) that align with the model’s behavior or "vibe." While this offers flexibility, it also makes the model’s behavior harder to predict. In contrast, we train five distinct LoRA adapters for four well-defined task types as defined in Table 2. These tasks are trained independently, with the base model’s weights kept frozen. For query-document retrieval, two adapters are trained jointly: one for queries and one for passages. During inference, users can select the appropriate adapter based on their downstream task and input role, ensuring optimal embeddings for their specific use case.
task value Task Description
retrieval.passage Embedding documents in a query-document retrieval task
retrieval.query Embedding queries in a query-document retrieval task
separation Clustering documents, visualizing a corpus
classification Text classification
text-matching Semantic text similarity, general symmetric retrieval, recommendation, finding similar items, deduplication
Table 2: Supported tasks of jina-embeddings-v3, each corresponding to a LoRA adapter and trained independently, except for retrieval.passage and retrieval.query, which are trained jointly.
@@ 4.3.1 Classification Adapter
The classification adapter generates embeddings that are effective for training downstream classification models, particularly logistic regression classifiers. To train the adapter, we employ the classification training method proposed for the Gecko embedding model Lee et al. (2024a). Specifically, we select datasets covering a range of common classification tasks, including sentiment analysis, intent classification, and article categorization.
@@
We employ an extended version of the InfoNCE loss $\mathcal{L}_{\text{triplet}}$ described in our previous work Günther et al. (2023) to take into account these additional negative samples.
@@ 4.3.4 Failure Analysis for Asymmetric Retrieval
The effectiveness of training the retrieval adapter on this data is evaluated in Section 5.4.
@@ 5.2 Performance on MTEB
Table 4 summarizes the performance of various multilingual text embedding models across different MTEB tasks, divided into English and multilingual sections. jina-embeddings-v3 performs competitively, particularly in monolingual English tasks, where it achieves the highest Classification Accuracy (CF) score of 82.58 and the top Sentence Similarity (STS) score of 85.80, demonstrating its robustness across both languages and tasks. Full evaluation results per task can be found in Appendix A3. When averaging across all tasks, jina-embeddings-v3 scores 65.52, outperforming models such as text-embedding-3-large, multilingual-e5-large-instruct, and Cohere-embed-multilingual-v3.0. This indicates that jina-embeddings-v3 maintains strong monolingual English performance while being trained on a wide variety of languages.
@@
For multilingual performance, Table 4 presents weighted averages across a wide selection of multilingual and cross-lingual MTEB tasks. The specific tasks, along with an explanation of which adapter was used for each task, are detailed in Appendix A7, A6, A8, A5, A9, and A4. Note that jina-embeddings-v2 refers to our bilingual model suite (jina-embeddings-v2-base-zh, jina-embeddings-v2-base-es, jina-embeddings-v2-base-de), which are evaluated only on Chinese, Spanish, and German monolingual and cross-lingual tasks, excluding all other tasks. Thus, the average score for jina-embeddings-v2* on the pair-classification tasks reflects performance on Chinese tasks alone. A complete report on pair-classification can be found in Appendix A6. We do not report scores for text-embedding-3-large and Cohere-embed-multilingual-v3.0 as these models were not evaluated on the full range of multilingual and cross-lingual MTEB tasks. However, our model outperforms multilingual-e5-large in all multilingual tasks except reranking and approaches the performance of multilingual-e5-large-instruct.